OpenStudy (anonymous):

How do you factor? For example 3x^2 - 12x^3=

OpenStudy (cwrw238):

take out the GCF whats GCF of the numbers 3 and 12?

OpenStudy (cwrw238):

GCF - greatest common factor

OpenStudy (anonymous):

and then?

OpenStudy (cwrw238):

then you find GCF of x^2 and x^3 and join it with first GCF

OpenStudy (cwrw238):

can u find first GCF - of 3 and 12?

OpenStudy (cwrw238):

can u find first GCF - of 3 and 12?

OpenStudy (cwrw238):

its the largest number that will divide into both 3 and 12

OpenStudy (anonymous):

the GCF is 3.

OpenStudy (anonymous):

and x the GCF is 3x

OpenStudy (cwrw238):

not quite

OpenStudy (anonymous):

because if you 3x by (x-4x^2) then you wud get ur orignial equation

OpenStudy (cwrw238):

GCF is 3x^2

OpenStudy (anonymous):

yup dats rite

OpenStudy (cwrw238):

and factors are 3x^2(1 - 4x)

OpenStudy (anonymous):

does this help u a liiltle more ?

OpenStudy (cwrw238):

x^3 / x^2 = x^(3-2) = x

OpenStudy (anonymous):

ok u lost me bt it al gud

OpenStudy (anonymous):

Thanks!

OpenStudy (cwrw238):

when you multiply things like x^3 , x^2 you add the powers x^3 * x^2 = x^(3+2) = x^5 when dividing u subtract the powers

OpenStudy (cwrw238):

yw lolkimkim

OpenStudy (anonymous):

u raise the powers

OpenStudy (anonymous):

ooohhh yeeah